# VocabTrim: Vocabulary Pruning for Efficient Speculative Decoding in LLMs

### v1

@@ 1 Introduction
Motivated by these observations, we propose VocabTrim, a training-free method for improving the efficiency of speculative decoding by reducing the size of the LM head of the drafter. To the best of our knowledge, this is the first approach that directly addresses the LM head, which has been largely overlooked in previous SpD research. By restricting the output dimension of the LM head, VocabTrim provides significant memory and latency savings without requiring any training or architectural changes. For Llama 3 models, we show that VocabTrim can reduce the output dimension of the LM heads by up to 75% with negligible impact to the acceptance rates. When applied to the state-of-the-art SpD method EAGLE-2, our method achieves on average 16% latency improvement on Spec-Bench tasks (Xia et al., 2024). We hope that this work encourages further research into optimizing previously underexplored components of the SpD pipeline.
@@ 2 Related Work
Recently, AdaptiVocab (Nakash et al., 2025) explored using custom vocabularies for LLM generation, to reduce computation and memory cost for target (or base) model. Note that the LM head linear layer consumes the most memory and computation during inference. In this work, we focus on LM head for drafter unlike Adaptivocab, to prevent any modification to target model’s final performance while exploiting the lossless property of SpD methods.

### v2

@@ 1 Introduction
Motivated by these observations, we propose VocabTrim, a training-free method for improving the efficiency of speculative decoding by reducing the size of the LM head of the drafter. By restricting the output dimension of the LM head, VocabTrim provides significant memory and latency savings without requiring any training or architectural changes. For Llama 3 models, we show that VocabTrim can reduce the output dimension of the LM heads by up to 75% with negligible impact to the acceptance rates. When applied to the state-of-the-art SpD method EAGLE-2, our method achieves on average 16% latency improvement on Spec-Bench tasks (Xia et al., 2024). We hope that this work encourages further research into optimizing previously underexplored components of the SpD pipeline.
@@ 2 Related Work
Recently, AdaptiVocab (Nakash et al., 2025) explored using custom vocabularies for LLM generation, to reduce computation and memory cost for target (or base) model. A concept similar to our approach was independently proposed in (Zhao et al., 2025), although their method is limited to pruning draft LM-head using only raw-text data. Detailed similarities and differences with (Zhao et al., 2025) are discussed in Section A.3. It is important to note that the LM head linear layer consumes a significant portion of memory and computation during inference. In contrast to AdaptiVocab, our work focuses specifically on optimizing the LM head of drafter. This design choice ensures that the final performance of the target model remains unaffected, while still leveraging the lossless nature of speculative decoding methods.
@@ 5 Conclusion
6 Acknowledgment
We thank Weilin Zhao and Xu Han who kindly pointed out the resemblance of the idea in our paper to their prior work (Zhao et al., 2025). We discuss their similarities and differences in Section 2 and Section A.3.
@@ A.2 Limitations
A.3 Comparison of VocabTrim and FR-Spec: Similarities and Differences
It has come to our attention that VocabTrim and FR-Spec (Zhao et al., 2025) share the idea of pruning LM head for speculative decoding. In this section, we highlight some key differences between them. First, they commonly identify the LM head as a bottleneck in draft model as motivation for the proposed works. While the induced algorithms resemble to each other, we emphasize that our focus is on the memory bandwidth of drafting process associated with the size of the drafter, regardless of its computational complexity. This is crucial in memory-bound environment, which is often the case in edge devices but not obvious in larger scale devices such as Nvidia GPUs in which cases the methods may lead to only marginal improvements. In this context, we report MBSU (memory-bound speed up) as the primary performance metric and make the performance measurement more consistent with on-edge-device performance.
Second, we propose the use of model-generated datasets, i.e., distillation datasets, to construct the token distribution, under the motivation of distribution matching and leading to the best performance on many cases, while comparing different types of calibration datasets: (a) target-generated, (b) draft-generated, and (c) raw dataset, similarly in (Zhao et al., 2025).
Lastly, our approach is a generic method for speculative decoding method with drafter LM head in the loop, not limited to EAGLE. Thus, we compare both independent drafter model and EAGLE-style drafter model.